OpenStudy (anonymous):

Express using rational exponents. CLICK 2 SEE.

OpenStudy (anonymous):

\[60 \sqrt[8]{r ^{80}s ^{56}t ^{27}}\]

OpenStudy (anonymous):

\[60(r^{80}s^{56}t^{27})^{\frac{ 1 }{ 8 }}\]

OpenStudy (anonymous):

could you explain how you go it ? like help me work it out ?

OpenStudy (anonymous):

\[\large \sqrt{2}=2^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

\[\large \sqrt[8]{2}=2^{\frac{ 1 }{ 8 }}\]

OpenStudy (anonymous):

Do you get it now?

OpenStudy (anonymous):

yes ! thanks :)

OpenStudy (anonymous):

You can expand the brackets now, right?

OpenStudy (anonymous):

you mean by distributing ?

OpenStudy (anonymous):

Yeah, if they call it that in your country.

OpenStudy (anonymous):

where do you live? lol ,

OpenStudy (anonymous):

Australia. We use the metric system and use terms that are different to what America uses and other countries that follow what the U.S does.

OpenStudy (anonymous):

that is very cool. but thanks for your help !

OpenStudy (anonymous):

My pleasure.